A balloon has a volume of 4.0 L when at sea level pressure of 760 mmHg at room temperature. What will be its volume when inflated at an elevation where the atmosphere pressure is 700 mmHg and at the same temperature.

1
Expert's answer
2021-12-21T04:16:02-0500

Since temperature is constant, we use Pressure law: P1V1 = P2V2

P1= 760mmHg, V1= 4.0L, P2= 700mmHg, V2=?

760*4 = 700*V2

3040 = 700*V2

V2= 3040/700 = 4.34L